Suppose an antidepressant medication has been shown to be 50% effective in treating depression. If the medication is given to ten unrelated individuals being treated for depression, what is the probability that it will be effective in more than 8? ( Round your answer to 4 decimal places)
Answer #1

X ~ bin ( n , p)

Where n = 10 , p = 0.5

Binomial probability distribution is

P(X) = nCx * px * ( 1 - p)n-x

P(X > 8) = P(X >= 9)

= P(X = 9) + P(X = 10)

= 10C9 * 0.59 * 0.5 + 10C10 * 0.510 * 0.50

= 0.0107